OpenStudy (tayidaniels):

Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%. Roger has 40 cars in his collection. Each year he collects 1 additional car. Part A: Write functions to represent Cory and Roger's collections throughout the years. (4 points) Part B: How many cars does Cory have after 6 years? How many does Roger have after the same number of years? (2 points) Part C: After approximately how many years is the number of cars that Cory and Roger have the same? Justify your answer mathematically. (4 point

OpenStudy (tayidaniels):

I already have Part A, but i don't know how to finish solving it. Part A: cory's: y=15(1.2)^x roger: y=x+40

OpenStudy (mathstudent55):

x represents the year, and y represents the number of cars in the collection. For part B, plug in 6 for x in both functions and find y.

OpenStudy (mathstudent55):

For part C: Cory's collection starts smaller but grows faster, so it will eventually catch up to Roger's collection. To find when it will catch up, do this. Let x = 1, and solve both functions for y. Then let x = 2, and solve both functions for y. Keep doing this with x = 3, x = 4, etc., until eventually, Cory's collection will be larger. The time in which Cory's collection catches up to Roger's collection is between the highest value of x that Cory's collection is still smaller and the lowest value of x in which Cory's collection is already larger. If you want to know the exact value of x, then set the two functions equal to each other, and solve for x.